Applied Managerial Statistics Homework: Week 3 Homework 40 pts Score: 0 of 5 pts 3.2.8 Save 60112 (4 complete) ▼ HW Score: 25%, 10 of 40 pts Question Help The table below shows the results of a survey in which 142 men and 145 women workers ages 25 to 64 were asked f they have at least one months income set aside for emergencies. Complete parts (a) through (d). Men Women Total Less than one months income 65 84 149 One months income or more 77 61 38 Total 142 145 287 (a) Find the probability that a randomly selected worker has one months income or more set aside for emergencies The probability is □ (Rou to the nearest thousandth as needed.) Enter your answer in the answer box and then click Check Answer Clear All Check Answer
Answer #1

P( One month's income or more) = n(One month's income or more) / n(Total)

= 138 / 287

= 0.481 (Rounded to nearest thousands)
Answer #2

SO.UTION :


P(worker has one month’s or more for emergencies)

= Possible outcomes / Total outcomes

= number of such workers / total number of workers

= 138/287 

= 0.481 (ANSWER)


answered by: Tulsiram Garg